# Multi-person eye tracking for real-world scene perception in social settings

### v1

@@ 4.1.6. Participants
A total of 60 people participated in the study (30 on each day, no repeated participants). Data from one participant was only recorded for half of the study due to dropout at intermission. Of the remaining 59 participants, 38 self-identified as women and the reported mean age was 34 years, ranging from 16 to 82 years. All participants had normal or corrected-to-normal vision (via contact lenses; participation while wearing another pair of glasses was not possible).
@@ 4.2.1. Time Synchronisation
Figure 5. A. Mean time offset (left) and mean roundtrip duration (right), in ms, recorded for the two sessions on each day. The coloured dots represent offsets at measured timepoints, (approx. 10 sec intervals); black dots represent mean values with the error bars representing standard error. B. Offset drift calculated for each device in each of the four sessions. The vertical bars represent standard deviation, horizontal bars represent the mean and individual dots represent single data points at measured timepoints.
@@
Overall, these data show relatively low offsets and drifts across all 30 devices and recording sessions. As is expected, mean offsets increase from the 1st to 2nd recording session for both groups (Fig. 5A; left panel), while drifts seem more so to be device-specific than recording session-specific. Regardless of the actual calculated offsets and drifts, our RANSAC approach allows us to accurately align the timestamps between each eye tracking mobile device and our central camera.
@@ 4.2.2. Visualisation
Figure 6. A. An example of feature matching from one frame of one participants’ worldview (left) onto the centralview (right). The recorded gaze corresponding to this frame is represented by the maroon and white circle. B. Multi-person views generated with the visualisation module. Top: All participants’ egocentric worldviews and gazes (small white and black circles) are displayed around the perimeter. Using the computed homography matrices for each participant, we project each participant’s gaze onto the centralview (centre). In the centre panel, participants’ gazes are represented with uniquely coloured circles with white in the middle to aid visibility. Bottom: Same as top, with the centre grid cell displaying a heatmap of the 2D gaze density of all participants looking at the scene. Higher intensities in the heatmap (intensity increases from blue to red with red being the highest) represent a higher proportion of the participants looking at that location. C. Same as centre panels in B, but resized to aid visibility.
@@
Raw egocentric gaze coordinates and homography-transformed gaze coordinates were projected onto each viewer’s worldview and the shared centralview, respectively, after removal of blinks. The visualisations (presented in Figure 6) provide a quick data check and demonstrate accurate projective transform from the SuperGlue features identified in the Homography module (see 3.2.5). The transforms handle video artefacts, such as motion blur (from sudden head movements), poor lighting/texture (frames with large dark regions, shadows, lens flare, glare etc.), obstructions (audience members in previous rows), and scale variations (different resolution and zoom in centralview and worldviews) reliably. However, further investigations and optimisations for edge cases in the scene where the projective transformation might fail could still be conducted and are out of the scope for the current paper. Figure 6 displays still, single frame examples of our transformed data; an example video can be viewed here: https://tinyurl.com/multipersonET. In the first part of this video example (tableau: “canjo”), the two performers are seated very near to each other. Even at this close distance, we can clearly distinguish gaze to one or the other performer, vs. their instruments, and follow shifts in audience attention between these elements. When this “canjo” tableau ends, the two performers spread out and we see the gaze follow them around the stage. Once one of the performers starts speaking (tableau: “pod rattle incident”); the audience mostly maintain their focus on him, rarely glancing toward the other performer who is gently striking metallic objects. This video highlights the observation that the transformed gaze reliably follows the two musicians around the stage and provides proof-of-concept that our homography approach successfully projects gaze from a large audience in challenging scenes.
@@
Heatmap analysis: Heatmaps help in analysing the 2D distribution of gaze over a time period. We present the individual gaze heatmaps of each participant for each of the four recording sessions in Figures 7 and 8, generated using the analysis module. Figure 7 shows the raw egocentric gaze data recorded from the eye tracking glasses in the worldview frame of reference. This egocentric gaze data is relative to the wearer’s head movements and demonstrates low variance in both horizontal and vertical directions with a strong central bias. This effect could be explained by the high coupling between eye and head movements commonly observed in real world exploration (Einhäuser et al., 2007; Tatler, 2007).
@@
To study the gaze patterns over an evolving scene and compare gaze data between different viewers it is, therefore, required to map gaze from the viewer’s frame of reference to the scene’s frame of reference. We achieve this with our Homography module (see 3.2.5) that maps each participant’s gaze to the shared centralview space. Figure 8 presents individual gaze heatmaps from all participants and the aggregate gaze activity from all participants for each session overlaid on an example frame from that session. Projecting the gaze to a central perspective gets rid of the independent positions and head movements of each participant, thereby getting rid of the central bias. Heatmaps of transformed gaze coordinates (Figure 8) demonstrate higher gaze dispersion and multiple discrete clusters for each session duration. These clusters coincide with the presentation screen area for the film viewing sessions and the performers’ movement activity for the concert performance sessions (see averaged heatmaps for each session in Figure 8). Higher gaze dispersion can be observed during performance viewing than during film viewing, on both days, which could be partially explained by the larger physical space covered on the stage by the performers during the concert, in comparison to the presentation size for the film (videowall screen).
@@
Investigating the gaze heatmaps for different participant seating clusters, we observe that participants in the right half of the audience spent more time exploring the right side of the stage, reflected by larger (higher gaze dispersion) and more intense (higher gaze duration) clusters in the right half and vice versa (see Figure 9A; heatmaps are arranged by seating order). This effect is, however, only observed in the performance sessions and not when the participants viewed the pre-recorded documentary film. The heatmaps could further be aggregated to observe similar effects with different participant clusters, such as clustering by participant columns (Figure 9A) or rows (Figure 9B). Figure 9 highlights different trends in the position of the heatmap centroid (marked by the intersection of white dashed lines; i.e., the peak heatmap intensity location), based on seating position.
@@
Figure 7. Egocentric gaze heatmaps. A. Participant-wise gaze heatmaps for each of the four event sessions. The heatmaps are arranged according to the seating order on both days where the top row (Row 1) is closest to the stage and the first column (from the left) is the leftmost participant. The sessions are arranged chronologically from top to bottom panel. B. Averaged gaze heatmaps calculated as the mean over all participants for each session.
@@
Figure 8. Transformed gaze heatmaps. A. Participant-wise gaze heatmaps for each of the event sessions. The heatmaps are arranged according to the seating order on both days where the top row (Row 1) is closest to the stage and the first column (from the left) is the leftmost participant. The sessions are arranged chronologically from top to bottom. B. Averaged gaze heatmaps calculated as the mean over all participants for each session.
@@ 4.2.3. Analysis
Spatial metrics: Further comparisons between individual heatmaps can be made with the help of metrics from the Analysis module (see 3.2.7). For example, pairwise SIM and CC scores were calculated for each participant pair for each of the four event sessions. For raw egocentric gaze, the mean SIM score during film viewing was 0.35 (SD = 0.21) for G1F and 0.38 (SD = 0.19) for G2F, while the mean SIM during the performance was 0.40 (SD = 0.18) for G1P and 0.43 (SD = 0.17) for G2P. The mean CC was 0.37 (SD = 0.30) for G1F, 0.42 (SD = 0.28) for G2P, 0.43 (SD = 0.26) for G1P and 0.46 (SD = 0.25) for G2P. When using homography-transformed gaze, the scores in all sessions, for both of these metrics, increase drastically (see Figure 10; left column), nearly doubling in most cases.
A 2D chart of seat positions (row, column) for each participant (see Figure 7 and 8 for audience seating positions on each day) was used to calculate euclidean distances (DI) between each participant pair. Calculated SIM scores on the homography-transformed gaze were found to be negatively correlated with the euclidean distances during the performance session on both days (i.e., higher SIM values for smaller seat distances between participants in a pair and vice versa). The Pearson correlation coefficient between SIM and DI was -0.10 (p = 0.039) for G1P and -0.17 (p ¡ 0.001) for G2P. These comparisons highlight the critical role of the homography transformation proposed in our framework for a rich analysis of multi-person eye tracking datasets. The trends and observations revealed by the transformed gaze data are not available from the raw egocentric gaze; gaze dispersion similarity between participants is much lower in the independent egocentric gaze coordinates and does not highlight clear patterns between participant pairs or groups.
Entropy and avg. velocity: The mean gaze entropy over all participants during film viewing was 0.76 (SD = 0.03) for G1F and 0.77 (SD = 0.03) for G2F. The mean entropy was higher when participants watched the performance, with a value of 0.82 (SD = 0.03) for G1P and 0.83 (SD = 0.03) for G2P. The higher values represent more spatial gaze dispersion when participants watch the performance as compared to the documentary film. The average gaze velocity (calculated in pixels/sample) during film viewing was found to be 22.99 (SD = 7.25) for G1P and 36.00 (SD = 5.90) for G2P. During the performance sessions, the average gaze velocity was 16.76 (SD = 6.79) for G1P and 22.11 (SD = 8.03) for G2P. Interestingly, gaze velocity was higher during film viewing on both days which signifies longer saccadic movements when viewing pre-recorded media on a screen, in comparison to a live concert performance
@@
Temporal metrics: A major contribution of our framework is the ability to study ocular activity over extended durations, as opposed to controlled short-duration trials in lab experimentation. It allows investigating long-term effects in complex tasks and capturing group dynamics in naturalistic social interactions. Leveraging the precise time synchronisation achieved with our framework, instantaneous ocular activity from multiple people can be investigated. For example, though we can report that the average blink duration for film is 302.93 (SD = 68.55) ms and performance is 306.49 (SD = 69.55) ms, it is much more interesting to be able to analyse blinks in a time resolved manner (Lange and Fink, 2023). Figure 11 demonstrates blinks from all 30 participants in each of the four recording sessions at discrete time points, over an example 1 minute duration. It highlights the variability of individual blink frequency and duration between participants in each session as well as the aggregate blink activity changes with time as each session progresses. Analysing such activity over the entire film or performance (each an hour+ in duration), will offer rich insights into collective information chunking, event segmentation, fatigue, immersion, etc.
@@
The synchronised gaze data from multiple devices mapped to a common coordinate space (centralview) further facilitates novel spatiotemporal investigations of collective gaze measures over time. In Figure 13, we visualise the measures of standard deviation, minimum convex area, and points in frame, over the entire duration of a film-viewing session (1 hr 20 min film, plus some minutes before/after) from 30 people (G1). These measures–computed by the Analysis module (see 3.2.7)–reveal collective gaze dynamics over time. Standard deviation of the gaze distribution of all participants at any given time is represented in the horizontal and vertical directions by the ”SD x” and ”SD y” measures (top and second panels, respectively). The ”Normalized convex area” (third panel) represents overall gaze dispersion as the area of a minimum convex hull enclosing gaze from all participants. The number of gaze points falling within the centralview scene at a given timepoint is represented by the ”Points in Frame” measure (bottom panel). The time series of all four measures highlight clear event boundaries such as the start and stop of the film presentation (vertical dotted lines). Future analyses will explore which features of the film drive changes in these collective gaze dynamics.
@@
Figure 13. A. Normalised contour area of the minimum convex hull enclosing all gaze points at each time point during the concert performance on both days (G1P: top; G2P: bottom). The performance consisted of 17 different tableaux; the dashed lines represent manually-annotated start times of each tableau in the performance sequentially with additional first and last lines representing start (performers entering stage) and end (end of last tableau) of the performance respectively. The colours correspond to the tableau number. All time series are smoothed with a rolling mean window size of 150 samples to filter high frequency noise for visualisation. B. Cross-Correlation matrix between each pair of tableaux. Green represents higher correlation between the normalised contour area metric during the two tableaux and red represents low correlation. C. Mean contour area of the time series segment for each tableau. The changes in the mean values are similar across the two days (G1P: Blue; G2P: Orange).
@@ 5. General Discussion
Results from our utility study demonstrate the end-to-end application of our framework in real-world social settings. The framework tackles 1) critical challenges of time-synchronisation of multi-sensor data to scale data collection, 2) monitoring and troubleshooting to ensure reliable data collection over long durations, 3) semantic gaze mapping of the synchronised gaze data to analyse collective gaze patterns, and finally, 4) shared gaze visualisation and analysis to achieve novel insights from the collected data. The proposed framework aims to reduce technological and logistical barriers in conducting naturalistic eye-tracking experiments with large social groups, allowing replication/refutation and extension of previous lab studies in social, naturalistic contexts.
The proposed method conceptualised and tackled key challenges related to multi-person eye-tracking, yet a major external dependency of our framework is the choice of eye-tracking hardware. We acknowledge that the availability/development of required functionality in existing eye trackers and affordability are still substantial barriers. We picked Pupil Labs Neon devices in our utility study over market competitors because of their remote API control, minimal calibration requirements, and the availability of open-source gaze-processing components. However, in our initial tests we encountered an array of operational issues on the proprietary vendor application that severely hindered reliable recording from the devices. Solving these issues required hours of testing, troubleshooting or downscaling. For example, the Neon devices are also capable of recording device orientation and movement at 110 Hz using the onboard 9-DoF Inertial Measurement Unit (IMU), as well as stereo audio from on-board microphones; we had to omit these data streams to allow reliable recording of the gaze and video data streams, which were our priority. IMU failures would crash the mobile application, even though the eye data were intact, and audio recording could not be reliably remotely started through API calls, as it requires the android device to be unlocked. These issues point out the very incipient stage of multi-person eye-tracking hardware/software since these systems are mostly tested for single person use cases. Moving forward, the technology would benefit from increased heavy testing, at the scale we propose in this paper. Further, the availability of open-source implementations would allow higher customizability for application specific requirements and thereby contribute to its adoption in novel settings.

### v2

@@ 4.1.6. Participants
A total of 60 people participated in the study (30 on each day, no repeated participants). Data from one of the participants was only recorded for half of the study due to dropout at intermission. Of the remaining 59 participants, 38 self-identified as women and the reported mean age was 34 years, ranging from 16 to 82 years. All participants had normal or corrected-to-normal vision (via contact lenses; participation while wearing another pair of glasses was not possible).
@@ 4.2.1. Time Synchronisation
    Figure 5. Time Synchronisation. A Mean time offset (left) and mean roundtrip duration (right), in ms, recorded for the two sessions on each day. The coloured dots represent offsets at measured timepoints, (approx. 10 sec intervals); black dots represent mean values with the error bars representing standard error. B Offset drift calculated for each device in each of the four sessions. The vertical bars represent standard deviation, horizontal bars represent the mean and individual dots represent single data points at measured timepoints.
@@
Overall, these data show relatively low offsets and drifts across all 30 devices and recording sessions. As is expected, mean offsets increase from the 1st to 2nd recording session for both groups (Fig. 5A; left panel), while drifts seem more so to be device-specific than recording session-specific. Regardless of the actual calculated offsets and drifts, our RANSAC approach allows us to accurately align the timestamps between each eye-tracking mobile device and our central camera.
@@ 4.2.2. Visualisation
Figure 6. Visualisations. A An example of feature matching from one frame of one participants’ worldview (left) onto the centralview (right). The recorded gaze corresponding to this frame is represented by the maroon and white circle. B Multi-person views generated with the visualisation module. Top: All participants’ egocentric worldviews and gazes (small white and black circles) are displayed around the perimeter. Using the computed homography matrices for each participant, we project each participant’s gaze onto the centralview (centre). In the centre panel, participants’ gazes are represented with uniquely coloured circles with white in the middle to aid visibility. Bottom: Same as top, with the centre grid cell displaying a heatmap of the 2D gaze density of all participants looking at the scene. Higher intensities in the heatmap (intensity increases from blue to red with red being the highest) represent a higher proportion of the participants looking at that location. C Same as centre panels in B, but resized to aid visibility.
@@
Raw egocentric gaze coordinates and homography-transformed gaze coordinates were projected onto each viewer’s worldview and the shared centralview, respectively, after removal of blinks. The visualisations (presented in Figure 6) enable data validation and demonstrate accurate projective transformation using the SuperGlue features identified in the Homography module (see 3.2.6). The transforms handle video artefacts, such as motion blur (from sudden head movements), poor lighting/texture (frames with large dark regions, shadows, lens flare, glare etc.), obstructions (audience members in previous rows), and scale variations (different resolution and zoom in centralview and worldviews) reliably. However, further optimisations and investigations would be needed to identify edge cases in the homography estimation. Figure 6 displays still, single frame examples of our transformed data; an example video can be viewed here: https://tinyurl.com/multipersonET. In the first part of this video example (tableau: “canjo”), the two performers are seated very near to each other. Even at this close distance, we can clearly distinguish audience-gaze on one or the other performer, vs. their instruments, and follow shifts in audience attention between these elements. When this “canjo” tableau ends, the two performers spread out and we see the gaze follow them around the stage. Once one of the performers starts speaking (tableau: “pod rattle incident”); the audience mostly maintain their focus on him, rarely glancing toward the other performer who is gently striking metallic objects. This video highlights the observation that the transformed gaze reliably follows the two musicians around the stage and provides proof-of-concept that our homography approach successfully projects gaze from a large audience in challenging scenes.
@@
We analyse the 2D distribution of gaze with the help of heatmaps (see Figures 7 and 8). Figure 7 shows the raw egocentric gaze data recorded from the eye-tracking glasses in the worldview coordinate space. This egocentric gaze data is relative to the wearer’s head movements and exhibits low variance in both horizontal and vertical directions with a strong central bias. This effect could be explained by the high coupling between eye and head movements commonly observed in real-world exploration (Einhäuser et al., 2007; Tatler, 2007).
@@
To study the gaze patterns over an evolving scene and compare gaze data between different viewers, we map gaze from viewer’s frame of reference to the scene’s frame of reference with our Homography module (see 3.2.6). Figure 8 presents individual gaze heatmaps from all participants and the aggregate gaze activity from all participants for each session overlaid on a reference frame from that session. Projecting the gaze to the static centerview perspective gets rid of the independent positions and head movements of each participant, thereby getting rid of the central bias. Heatmaps of transformed gaze coordinates (Figure 8) demonstrate higher gaze dispersion and multiple discrete clusters for each session duration. These clusters coincide with the presentation screen area for the film viewing sessions and the performers’ movement activity for the concert performance sessions (see averaged heatmaps for each session in Figure 8). Higher gaze dispersion can be observed during performance viewing than during film viewing, on both days, which could be partially explained by the larger physical space covered on the stage by the performers during the concert, in comparison to the presentation screen-size for the film.
@@
Investigating the gaze heatmaps for different participant seating clusters, we observe that participants in the right half of the audience spent more time exploring the right side of the stage, reflected by larger (higher gaze dispersion) and more intense (higher gaze duration) clusters in the right half and vice versa (see Figure 9A; heatmaps are arranged by seating order). This effect is, however, only observed in the performance sessions and not when the participants viewed the pre-recorded documentary film. The heatmaps could further be aggregated to observe similar effects with different participant clusters, such as clustering by participant columns (Figure 9A) or rows (Figure 9B). Figure 9 highlights different trends in the position of the heatmap centroid (i.e., the peak heatmap intensity location; marked by the intersection of white dashed lines), based on seating position.
@@
Figure 7. Egocentric gaze heatmaps. A Participant-wise gaze heatmaps for each of the four event sessions. The heatmaps are arranged according to the seating order on both days where the top row (Row 1) is closest to the stage and the first column (from the left) is the leftmost participant. The sessions are arranged chronologically from top to bottom panel. B Averaged gaze heatmaps calculated as the mean over all participants for each session.
@@
Figure 8. Transformed gaze heatmaps. A Participant-wise gaze heatmaps for each of the event sessions. The heatmaps are arranged according to the seating order on both days where the top row (Row 1) is closest to the stage and the first column (from the left) is the leftmost participant. The sessions are arranged chronologically from top to bottom. B Averaged gaze heatmaps calculated as the mean over all participants for each session.
@@ 4.2.3. Analysis
Further comparisons between individual heatmaps can be made with the help of metrics from our Analysis module (see 3.2.8). Pairwise SIM and CC scores were calculated for each participant pair for each of the four event sessions. For raw egocentric gaze, the mean SIM score during film viewing was 0.35 (SD = 0.21) for G1F and 0.38 (SD = 0.19) for G2F, while the mean SIM during the performance was 0.40 (SD = 0.18) for G1P and 0.43 (SD = 0.17) for G2P. The mean CC was 0.37 (SD = 0.30) for G1F, 0.42 (SD = 0.28) for G2P, 0.43 (SD = 0.26) for G1P and 0.46 (SD = 0.25) for G2P. When using homography-transformed gaze, the scores in all sessions, for both of these metrics, increase drastically (see Figure 10; left column), nearly doubling in most cases. A 2D chart of seat positions (row, column) for each participant (see Figure 7 and 8 for audience seating positions on each day) was used to calculate euclidean distances (DI) between each participant pair. Calculated SIM scores on the homography-transformed gaze were found to be negatively correlated with the euclidean distances during the performance session on both days (i.e., higher SIM values for smaller seat distances between participants in a pair and vice versa). The Pearson correlation coefficient between SIM and DI was -0.10 (p = 0.039) for G1P and -0.17 (p ¡ 0.001) for G2P.
These comparisons highlight the critical role of the homography transformation proposed in our framework for a rich analysis of multi-person eye-tracking datasets. The trends and observations revealed by the transformed gaze data are not available from the raw egocentric gaze; gaze dispersion similarity between participants is much lower in the independent egocentric gaze coordinates and does not highlight clear patterns between participant pairs or groups.
The mean gaze entropy over all participants during film viewing was 0.76 (SD = 0.03) for G1F and 0.77 (SD = 0.03) for G2F. The mean entropy was higher when participants watched the performance, with a value of 0.82 (SD = 0.03) for G1P and 0.83 (SD = 0.03) for G2P. The higher values represent more spatial gaze dispersion when participants watch the performance as compared to the documentary film. The average gaze velocity (calculated in pixels/sample) during film viewing was found to be 22.99 (SD = 7.25) for G1P and 36.00 (SD = 5.90) for G2P. During the performance sessions, the average gaze velocity was 16.76 (SD = 6.79) for G1P and 22.11 (SD = 8.03) for G2P. Interestingly, gaze velocity was higher during film viewing on both days which signifies longer saccadic movements when viewing pre-recorded media on a screen, in comparison to a live concert performance
@@
4.2.4. Exploratory Temporal Analysis
A major contribution of SocialEyes is the ability to study ocular activity over extended durations, as opposed to controlled short-duration trials in lab experimentation. It allows investigating long-term effects in complex tasks and capturing group dynamics in naturalistic social interactions. Leveraging the precise time synchronisation achieved with our framework, instantaneous ocular activity from multiple people can be investigated. For example, though we can report that the average blink duration for film is 302.93 (SD = 68.55) ms and performance is 306.49 (SD = 69.55) ms, it is much more interesting to be able to analyse blinks in a time resolved manner (Lange and Fink, 2023). Figure 11 demonstrates blinks from all 30 participants in each of the four recording sessions at discrete time points, over a randomly-clipped 1 minute section. It highlights the variability of individual blink frequency and duration between participants in each session as well as the aggregate blink activity changes with time as each session progresses. Analysing such activity over the entire film or performance (each an hour+ in duration), will offer rich insights into collective information chunking, event segmentation, fatigue, immersion, etc.
@@ 4.2.4. Exploratory Temporal Analysis
The synchronised gaze data from multiple devices mapped to a common coordinate space (centralview) further facilitates novel spatiotemporal investigations of collective gaze measures over time. In Figure 12, we visualise the measures of standard deviation, minimum convex area, and points in frame, over the entire duration of a film-viewing session (1 hr 20 min film, plus some minutes before/after) from 30 people (G1). These measures–computed by the Analysis module (see 3.2.8)–reveal collective gaze dynamics over time. Standard deviation of the gaze distribution of all participants at any given time is represented in the horizontal and vertical directions by the ”SD x” and ”SD y” measures (top and second panels, respectively). The ”Normalized convex area” (third panel) represents overall gaze dispersion as the area of a minimum convex hull enclosing gaze from all participants. The number of gaze points falling within the centralview scene at a given timepoint is represented by the ”Points in Frame” measure (bottom panel). The time series of all four measures highlight clear event boundaries such as the start and stop of the film presentation (vertical dotted lines). Future analyses will explore which features of the film drive changes in these collective gaze dynamics.
@@
Figure 13. Gaze behaviour during concert viewing. A Normalised contour area of the minimum convex hull enclosing all gaze points at each time point during the concert performance on both days (G1P: top; G2P: bottom). The performance consisted of 17 different tableaux; the dashed lines represent manually-annotated start times of each tableau in the performance sequentially with additional first and last lines representing start (performers entering stage) and end (end of last tableau) of the performance respectively. The colours correspond to the tableau number. All time series are smoothed with a rolling mean window size of 150 samples to filter high frequency noise for visualisation. B Cross-Correlation matrix between each pair of tableaux. Green represents higher correlation between the normalised contour area metric during the two tableaux and red represents low correlation. C Mean contour area of the time series segment for each tableau. The changes in the mean values are similar across the two days (G1P: Blue; G2P: Orange).
@@ 5. General Discussion
Results from our utility test demonstrate the end-to-end implementation of the recording mode of SocialEyes in a real-world application. SocialEyes tackles 1) critical challenges of time-synchronisation of multi-sensor data to scale data collection, 2) monitoring and troubleshooting to ensure reliable data collection over long durations, 3) semantic gaze mapping of the synchronised gaze data to analyse collective gaze patterns, and finally, 4) shared gaze visualisation and analysis to achieve novel insights from the collected data. The proposed framework aims to reduce technological and logistical barriers in conducting naturalistic eye-tracking experiments with large social groups, allowing replication/refutation and extension of previous lab studies in social, naturalistic contexts (Kulke and Hinrichs, 2021; Laidlaw et al., 2016; Foulsham and Kingstone, 2017).
The proposed method conceptualised and tackled key challenges related to multi-person eye-tracking, yet a major external dependency of our framework is the choice of eye-tracking hardware. We acknowledge that the availability/development of required functionality in existing eye trackers and affordability are still substantial barriers. We picked Pupil Labs Neon devices in our utility test over market competitors because of their remote API control, minimal calibration requirements, and the availability of open-source gaze-processing components. However, in our initial tests we encountered an array of operational issues on the proprietary vendor application that severely hindered reliable recording from the devices. Solving these issues required hours of testing, troubleshooting or downscaling. For example, the Neon devices are also capable of recording device orientation and movement at 110 Hz using the onboard 9-DoF Inertial Measurement Unit (IMU), as well as stereo audio from on-board microphones; we had to omit these data streams to allow reliable recording of the gaze and video data streams, which were our priority. IMU failures would crash the mobile application, even though the eye data were intact, and audio recording could not be reliably remotely started through API calls, as it requires the android device to be unlocked. These issues point out the very incipient stage of multi-person eye-tracking hardware/software since these systems are mostly tested for single person use cases. Moving forward, the technology would benefit from increased heavy testing, at the scale we propose in this paper. Further, the availability of open-source implementations would allow higher customizability for application specific requirements and thereby contribute to its adoption in novel settings.